OpenStudy (josee):

is there a number that is exactly 1 more than its cube? if so what what? and how do you do this..lol

OpenStudy (anonymous):

just starting the problem, not sure of the answer yet. the way to set up this problem is like this: x^3=x-1

OpenStudy (anonymous):

and actually that's impossible. just decided.

OpenStudy (josee):

it is possible. lol

OpenStudy (josee):

the book is saying there is a number. but that's all it says

OpenStudy (anonymous):

the reason being: 1^3 = 1 so anything less than 1 is going to have a cube smaller than itself. but if you subtract one from any of those numbers you will get a negative number which is less than 1 away from its cube

OpenStudy (anonymous):

could you restate the question? you are looking for a number that is exactly one more than it's cube?

OpenStudy (josee):

i looked around and it says that -1.32...is the number

OpenStudy (josee):

yea that's the question

OpenStudy (anonymous):

your answer is right according to wolfram, but i don't know how they got it, sorry bro

OpenStudy (josee):

:/